LowkeyPanda:

I really need help because im so bad at algebra Choose the correct simplification of the expression the quantity a to the 4th power over b to the 2nd power all raised to the 3rd power . (1 point) a to the 7th power over b to the 5th power a to the 12th power over b to the 6th power ab9 a12b6

qwertyiop:

First off what do you think it is?

LowkeyPanda:

i dont know because i'm bad at it that why i need help

LowkeyPanda:

sorry i had to grab my step dads wallet

qwertyiop:

Have you tried solving it yet?

LowkeyPanda:

i don't really know how

LowkeyPanda:

also sorry for any inconvenience

qwertyiop:

I have to go sorry ): @supie

LowkeyPanda:

thank you for trying to help me

qwertyiop:

I feel really bad tho

jhonyy9:

a to the 4th power over b to the 2nd power all raised to the 3rd power . (1 point)\[\left( \frac{ a^{4} }{ b^{2} } \right)^3 = \frac{ (a^{4})^3 }{ (b^{2})^3 } = ? \] any idea now ? @XioGonz